# Teaching LLMs How to Learn with Contextual Fine-Tuning

### v1

@@ Results.
Figure 3: Financial Benchmarks (Zero-shot). This table presents the impact of negative contextual prompts on financial benchmarks. It shows a notable performance drop for the 7B model, with a decrease of %$\%\Delta^{-CFT}_{CFT}=-3.41\%$, and a smaller yet significant reduction for the 13B model at $-2.39\%$.
@@ Appendix E Additional Experiments
”I have the following medical question from MedQA USMLE exam to prepare for: Q: {QUESTION} A: {OPTION A} B: {OPTION B} C: {OPTION C} D: {OPTION D} Answer: {CORRECT OPTION}. {ANSWER} Please provide: A thorough textbook-style explanation written in clear, connected paragraphs that build upon each other. Include relevant clinical correlations and physiological mechanisms throughout the text. Present this as a cohesive educational yet concise resource similar to what I might find in a high-quality medical textbook structure. The output should have less than 3000 characters.”
@@
”Based on the following question-answer pair and its related educational content: QUESTION: {QUESTION} A: {OPTION A} B: {OPTION B} C: {OPTION C} D: {OPTION D} Answer: {CORRECT OPTION}. {ANSWER} EDUCATIONAL CONTENT: {EDUCATIONAL CONTENT} Generate a very concise contextual prompt that will enhance learning effectiveness. The prompt should: 1. Follow the style of [select one learning theory approach: Application of Knowledge/In-Depth Exploration/Reflective Thinking/Creative Interpretation/Summarization and Synthesis/Focus on Key Concepts/Contextual Understanding/Critical Analysis/Question-Based Learning/Comparative Learning] 2. Identify: The fundamental concepts that must be understood Critical facts that require focus for mastery How these elements connect to clinical reasoning or application 3. Be formatted as a directive that encourages active engagement with the material (approximately 1-2 sentences) 4. Frame the learning in a way that facilitates long-term retention The contextual prompt should help the learner not just memorize information but develop a deeper, more applicable understanding of the medical concept to correctly answer the question using the educational content. The output should only contain the concise contextual prompt with 1-2 sentences.”
@@ E.2 Baseline Comparison
We compare CFT with AdaptLLM (Cheng et al., 2024), a domain-specific continued pretraining method that enhances knowledge by converting corpora into a reading comprehension format for fine-tuning. As shown in Table 4, CFT consistently outperforms AdaptLLM across all tasks in the biomedical benchmarks. In contrast, AdaptLLM generally underperforms on our dataset. We attribute this discrepancy to differences in the datasets used. he original AdaptLLM paper utilizes PubMed abstracts to create reading comprehension tasks. Abstracts typically provide concise summaries of articles, making it easier to generate meaningful question-answer pairs. In contrast, our dataset comprises full-text articles from MDPI journals and textbooks, where not every paragraph is suitable for question-answer generation. This limitation may reduce the effectiveness of AdaptLLM’s approach on our data. This suggests that our method is more efficient, simpler, and data-structure-agnostic compared to AdaptLLM.
@@ E.3 Evaluation on General and Instruction-Following Benchmarks
We evaluate the OpenMedText fine-tuned Llama-2-13B model on general benchmarks and instruction-following benchmarks: (1) Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021), (2) MMLU-Pro (Wang et al., 2024), and (3) Instruction-Following Eval (IFEval) (Zhou et al., 2023). As shown in Table 13, fine-tuning does not cause catastrophic forgetting; the model’s general knowledge is only slightly diminished, and the capabilities of the base model are largely retained. While CPT is marginally more robust to knowledge degradation than CFT, the performance difference is minimal and CFT demonstrates stronger in-domain performance, as evidenced by Table 1.
@@ Appendix F Understanding Contextual fine-tuning with Synthetic Data
Briefly, Garg et al. (2022) aim to train a transformer model capable of in-context learning of a function class $\mathcal{F}$. Their goal is to demonstrate that after sufficient training, transformers can approximate any function $f\in\mathcal{F}$ by conditioning on a sequence of in-context examples at inference time. Building upon their framework, we adopt their pretraining setup to train a model that learns a class of function $\mathcal{F}$. However, our objective differs in that we investigate how contextual fine-tuning affects the pretrained model’s ability to learn a new function class $\mathcal{G}$. Specifically, consider a function class $\mathcal{F}$, our initial goal is to train a model that can learn functions $f\in\mathcal{F}$ such that, for most functions, the model can approximate $f(x_{\text{query}})$ for a new query input $x_{\text{query}}$ by conditioning on a prompt sequence containing in-context examples. Formally, let $\mathcal{D}_{\mathcal{X}}$ be a distribution over inputs, and let $\mathcal{D}_{\mathcal{F}}$ be a distribution over functions in $\mathcal{F}$.
@@
where $\ell$ is the mean squared error loss. In our experiment, $\mathcal{F}$ is the class of linear functions, that is, =$\mathcal{F}=\{f\mid f(x)=w^{\top}x,w\in\mathbb{R}^{d}\}$, where the weight vectors $w$ are sampled from $\mathcal{N}(0,I_{d})$. We let $\mathcal{D}_{\mathcal{X}}$ be the isotropic Gaussian distribution $\mathcal{N}(0,I_{d})$. Garg et al. (2022) show that after sufficient training, a transformer model can predict $f(x_{\text{query}})$ almost perfectly when there are more than 20 in-context examples.
@@ Results.
Contextual fine-tuning improves learning dynamics. Figure 4(a) and 4(c) illustrate that transformers fine-tuned using CFT achieve lower loss compared to those trained with CPT and NEG-CFT suggesting that the content of the contextual prompts in both cases better guides training dynamics when learning a new function class $\mathcal{G}$. In Figure 4(b) and 4(d), we assess the model’s performance at test time with normalized squared error $(M_{\theta}(x)-g(x))^{2}/d$ where $d=20$ is the dimensionality of the input and weight vectors. The contextual fine-tuned transformer achieves lower errors even with a small number of in-context examples in both the polynomial combination and multiple linear relationships case. This demonstrates that CFT helps the model to learn the function class $\mathcal{G}$ more accurately than existing training strategies.
@@
Clotting proteins are mainly produced in the liver. Twelve proteins known as “clotting factors” participate in the cascade clotting process during endothelial injury. One important clotting factor is fibrinogen. Fibrinogen generates fibrin when activated by the coagulant thrombin, which forms a mesh that clots blood with the assistance of a platelet plug. Normally, anticoagulants and fibrinolytics in the plasma, such as plasmin and heparin, break up fibrin clots and inactivate thrombin. However, during endothelial injury, damaged cells will release tissue factor, another type of clotting factor that causes a cascade of thrombin production that will overpower the anticoagulants and cause a clotting response. Serum is a term used to describe plasma that has been removed of its clotting factors. Serum still contains albumin and globulins, which are often called serum proteins as a result. Curation and Revision. Authored by: Boundless.com. Provided by: Boundless.com. License: CC BY-SA: Attribution- ShareAlike Blood. Provided by: Wikipedia. Located at: en.Wikipedia.org/wiki/Blood%2…of_human_blood. License: CC BY-SA: Attribution-ShareAlike Human Physiology/Blood physiology. Provided by: Wikibooks. Located at: en.wikibooks.org/wiki/Human_Composition. License: CC BY-SA: Attribution-ShareAlike plasma. Provided by: Wiktionary. Located at: en.wiktionary.org/wiki/plasma. License: CC BY-SA: Attribution-ShareAlike glycoprotein. Provided by: Wiktionary. Located at: en.wiktionary.org/wiki/glycoprotein.
@@
Clotting proteins are primarily produced in the liver. Twelve proteins, known as "clotting factors," participate in the cascade clotting process during endothelial injury. One crucial clotting factor is fibrinogen. Fibrinogen generates fibrin when activated by thrombin, which forms a mesh that clots blood with the assistance of a platelet plug. Anticoagulants and fibrinolytics in the plasma, such as plasmin and heparin, normally break up fibrin clots and inactivate thrombin. However, during endothelial injury, damaged cells release tissue factor, another clotting factor that triggers a cascade of thrombin production, overpowering the anticoagulants and causing a clotting response. Serum refers to plasma without clotting factors. Serum still contains albumin and globulins, which are often referred to as serum proteins.

### v2

@@ Results.
Figure 3: Financial Benchmarks (Zero-shot).
@@ Appendix E Additional Experiments
“I have the following medical question from MedQA USMLE exam to prepare for: Q: {QUESTION} A: {OPTION A} B: {OPTION B} C: {OPTION C} D: {OPTION D} Answer: {CORRECT OPTION}. {ANSWER} Please provide: A thorough textbook-style explanation written in clear, connected paragraphs that build upon each other. Include relevant clinical correlations and physiological mechanisms throughout the text. Present this as a cohesive educational yet concise resource similar to what I might find in a high-quality medical textbook structure. The output should have less than 3000 characters.”
@@
“Based on the following question-answer pair and its related educational content: QUESTION: {QUESTION} A: {OPTION A} B: {OPTION B} C: {OPTION C} D: {OPTION D} Answer: {CORRECT OPTION}. {ANSWER} EDUCATIONAL CONTENT: {EDUCATIONAL CONTENT} Generate a very concise contextual prompt that will enhance learning effectiveness. The prompt should: 1. Follow the style of [select one learning theory approach: Application of Knowledge/In-Depth Exploration/Reflective Thinking/Creative Interpretation/Summarization and Synthesis/Focus on Key Concepts/Contextual Understanding/Critical Analysis/Question-Based Learning/Comparative Learning] 2. Identify: The fundamental concepts that must be understood Critical facts that require focus for mastery How these elements connect to clinical reasoning or application 3. Be formatted as a directive that encourages active engagement with the material (approximately 1-2 sentences) 4. Frame the learning in a way that facilitates long-term retention The contextual prompt should help the learner not just memorize information but develop a deeper, more applicable understanding of the medical concept to correctly answer the question using the educational content. The output should only contain the concise contextual prompt with 1-2 sentences.”
@@ E.2 Baseline Comparison
We compare CFT with AdaptLLM (Cheng et al., 2024), a domain-specific continued pretraining method that enhances knowledge by converting corpora into a reading comprehension format for fine-tuning. As shown in Table 4, CFT consistently outperforms AdaptLLM across all tasks in the biomedical benchmarks. In contrast, AdaptLLM generally underperforms on our dataset. We attribute this discrepancy to differences in the datasets used. The original AdaptLLM paper utilizes PubMed abstracts to create reading comprehension tasks. Abstracts typically provide concise summaries of articles, making it easier to generate meaningful question-answer pairs. In contrast, our dataset comprises full-text articles from MDPI journals and textbooks, where not every paragraph is suitable for question-answer generation. This limitation may reduce the effectiveness of AdaptLLM’s approach on our data. This suggests that our method is more efficient, simpler, and data-structure-agnostic compared to AdaptLLM.
@@ E.3 Evaluation on General and Instruction-Following Benchmarks
We evaluate the OpenMedText fine-tuned Llama-2-13B model on general benchmarks and instruction-following benchmarks: (1) Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021), (2) MMLU-Pro (Wang et al., 2024), and (3) Instruction-Following Eval (IFEval) (Zhou et al., 2023). As shown in Table 13, fine-tuning does not cause catastrophic forgetting; the model’s general knowledge is only slightly diminished, and the capabilities of the base model are largely retained. While CPT is marginally more robust to knowledge degradation than CFT, the performance difference is minimal, and CFT demonstrates stronger in-domain performance, as evidenced by Table 1.
@@ Appendix F Understanding Contextual fine-tuning with Synthetic Data
Briefly, Garg et al. (2022) aim to train a transformer model capable of in-context learning of a function class $\mathcal{F}$. Their goal is to demonstrate that after sufficient training, transformers can approximate any function $f\in\mathcal{F}$ by conditioning on a sequence of in-context examples at inference time. Building upon their framework, we adopt their pretraining setup to train a model that learns a class of functions $\mathcal{F}$. However, our objective differs in that we investigate how contextual fine-tuning affects the pretrained model’s ability to learn a new function class $\mathcal{G}$. Specifically, consider a function class $\mathcal{F}$, our initial goal is to train a model that can learn functions $f\in\mathcal{F}$ such that, for most functions, the model can approximate $f(x_{\text{query}})$ for a new query input $x_{\text{query}}$ by conditioning on a prompt sequence containing in-context examples. Formally, let $\mathcal{D}_{\mathcal{X}}$ be a distribution over inputs, and let $\mathcal{D}_{\mathcal{F}}$ be a distribution over functions in $\mathcal{F}$.
@@
where $\ell$ is the mean squared error loss. In our experiment, $\mathcal{F}$ is the class of linear functions, that is, $\mathcal{F}=\{f\mid f(x)=w^{\top}x,w\in\mathbb{R}^{d}\}$, where the weight vectors $w$ are sampled from $\mathcal{N}(0,I_{d})$. We let $\mathcal{D}_{\mathcal{X}}$ be the isotropic Gaussian distribution $\mathcal{N}(0,I_{d})$. Garg et al. (2022) show that after sufficient training, a transformer model can predict $f(x_{\text{query}})$ almost perfectly when there are more than 20 in-context examples.
@@ Results.
Contextual fine-tuning improves learning dynamics. Figure 4(a) and 4(c) illustrate that transformers fine-tuned using CFT achieve lower loss compared to those trained with CPT and NEG-CFT, suggesting that the content of the contextual prompts in both cases better guides training dynamics when learning a new function class $\mathcal{G}$. In Figure 4(b) and 4(d), we assess the model’s performance at test time with normalized squared error $(M_{\theta}(x)-g(x))^{2}/d$ where $d=20$ is the dimensionality of the input and weight vectors. The contextual fine-tuned transformer achieves lower errors even with a small number of in-context examples in both the polynomial combination and multiple linear relationships cases. This demonstrates that CFT helps the model to learn the function class $\mathcal{G}$ more accurately than existing training strategies.
@@
Clotting proteins are mainly produced in the liver. Twelve proteins known as “clotting factors” participate in the cascade clotting process during endothelial injury. One important clotting factor is fibrinogen. Fibrinogen generates fibrin when activated by the coagulant thrombin, which forms a mesh that clots blood with the assistance of a platelet plug. Normally, anticoagulants and fibrinolytics in the plasma, such as plasmin and heparin, break up fibrin clots and inactivate thrombin. However, during endothelial injury, damaged cells will release tissue factor, another type of clotting factor that causes a cascade of thrombin production that will overpower the anticoagulants and cause a clotting response. Serum is a term used to describe plasma that has been removed of its clotting factors. Serum still contains albumin and globulins, which are often called serum proteins as a result. Curation and Revision. Authored by: Boundless.com. Provided by: Boundless.com. License: $C$ BY-SA: Attribution- ShareAlike Blood. Provided by: Wikipedia. Located at: en.Wikipedia.org/wiki/Blood%2…of_human_blood. License: $C$ BY-SA: Attribution-ShareAlike Human Physiology/Blood physiology. Provided by: Wikibooks. Located at: en.wikibooks.org/wiki/Human_Composition. License: $C$ BY-SA: Attribution-ShareAlike plasma. Provided by: Wiktionary. Located at: en.wiktionary.org/wiki/plasma. License: $C$ BY-SA: Attribution-ShareAlike glycoprotein. Provided by: Wiktionary. Located at: en.wiktionary.org/wiki/glycoprotein.
@@
Clotting proteins are primarily produced in the liver. Twelve proteins, known as ‘‘clotting factors,’’ participate in the cascade clotting process during endothelial injury. One crucial clotting factor is fibrinogen. Fibrinogen generates fibrin when activated by thrombin, which forms a mesh that clots blood with the assistance of a platelet plug. Anticoagulants and fibrinolytics in the plasma, such as plasmin and heparin, normally break up fibrin clots and inactivate thrombin. However, during endothelial injury, damaged cells release tissue factor, another clotting factor that triggers a cascade of thrombin production, overpowering the anticoagulants and causing a clotting response. Serum refers to plasma without clotting factors. Serum still contains albumin and globulins, which are often referred to as serum proteins.